# A Robust Real-Time Lane Detection Method with Fog-Enhanced Feature Fusion for Foggy Conditions

### v10

@@ Abstract
Lane detection is a critical component of Advanced Driver Assistance Systems (ADAS). Existing lane detection algorithms generally perform well under favorable weather conditions. However, their performance degrades significantly in adverse conditions, such as fog, which increases the risk of traffic accidents. This challenge is compounded by the lack of specialized datasets and methods designed for foggy environments. To address this, we introduce the FoggyLane dataset, a diverse and comprehensive benchmark captured in real-world foggy scenarios across multiple road types and conditions, and synthesize two more datasets, FoggyCULane and FoggyTusimple, derived from widely used lane detection datasets. Furthermore, we propose a robust Fog-Enhanced Network for lane detection, incorporating a Global Feature Fusion Module (GFFM) to capture global relationships in foggy images, a Kernel Feature Fusion Module (KFFM) to model the structural and positional relationships of lane instances, and a Low-level Edge Enhanced Module (LEEM) to address missing edge details in foggy conditions. Comprehensive experiments demonstrate that our method achieves state-of-the-art performance, with F1-scores of 95.04 on FoggyLane, 79.85 on FoggyCULane, and 96.95 on FoggyTusimple. Additionally, with TensorRT acceleration, the method reaches a processing speed of 38.4 FPS on the NVIDIA Jetson AGX Orin, confirming its real-time capabilities and robustness in foggy environments.
@@
ADAS, Intelligent vehicle, Foggy lane detection dataset, Foggy lane detection, Global attention, Edge enhancement
journal: Neurocomputing
@@
[1]organization=Guangdong Key Laboratory of Intelligent Transportation System, School of Intelligent Systems Engineering, Sun Yat-sen University, city=Guangzhou, postcode=510275, country=China
[…]
[2]organization=College of Automotive Studies, Tongji University, city=Shanghai, postcode=201804, country=China
[…]
[3]organization=College of Transportation Engineering, Tongji University, city=Shanghai, postcode=201804, country=China
[…]
[4]organization=Department of Civil and Environmental Engineering, University of Alberta, city=Edmonton, state=Alberta, country=Canada
[…]
[5]organization=State Key Laboratory of Automotive Chassis Integration and Bionics, Jilin University, city=Changchun, postcode=130025, country=China
{highlights}
A new real-world dataset and two synthetic datasets for foggy lane detection are proposed.
A novel Fog-Enhanced Network is proposed for lane detection in fog.
Three core modules are designed to enhance feature extraction in fog.
The method achieves SOTA performance with real-time speed on embedded devices.
@@ 1 Introduction
Lane detection plays an important role in Advanced Driver Assistance Systems (ADAS), providing essential support for lane keeping, autonomous navigation, and lane departure warnings, which help reduce the risk of lane-deviation accidents[1]. As intelligent driving technologies advance, the need for robust and real-time lane detection in complex road environments is becoming more apparent, making it a key research focus in autonomous driving[2].
@@
Under good weather and clear lane markings, current lane detection algorithms perform well. However, in complex road environments and poor weather, such as those with poor weather conditions like fog, lane markings can become blurred, image quality may decrease, and global environmental information is often obscured, which makes lane detection more challenging. In these conditions, existing algorithms often struggle and may fail, leading to an increased risk of traffic accidents with more severe consequences. Ensuring safe driving in these conditions calls for the development of high-precision, highly robust lane detection methods optimized for foggy environments. As shown in Fig. 1, the in-vehicle camera captures real-time road images, which are preprocessed by edge devices and then sent to the lane detection module. Detected lane information is fused with data from other sensors, such as radar and GPS, along with real-time traffic data from the cloud. This information is processed in the vehicle’s computing unit, allowing the system to make adaptive driving decisions—such as accelerating, braking, or steering—to ensure safety in challenging environments like fog.
Figure 1: Lane detection in foggy scenarios for ADAS.[3]
@@
Earlier lane detection algorithms [4, 5] primarily relied on traditional computer vision methods, such as feature-based approaches that detect vanishing points, use color features, and apply edge detection or line detection techniques like Hough transforms. These methods are efficient for real-time applications but require manual parameter tuning and often struggle in challenging scenarios like worn or occluded lane markings.
@@
With the rise of deep learning, many modern lane detection methods [6, 7, 8, 9, 10, 11] leverage CNNs, which automatically extract features from raw data and generalize better in complex road environments. While deep learning provides higher robustness, many models still rely on pixel-by-pixel classification, which is computationally expensive and unnecessary for lane detection. Moreover, lane detection doesn’t always require dense pixel segmentation, which can be inefficient.
@@
Additionally, some approaches have incorporated anchor-based methods from object detection to improve efficiency and accuracy, particularly for complex or occluded lanes. These methods preset anchors at different starting points and slopes to optimize lane shape and position prediction [12, 13, 14, 15, 16]. However, these techniques still face challenges in handling foggy conditions, where image degradation severely impacts lane visibility. Traditional image enhancement methods, like dehazing, are commonly used to improve clarity but do not always result in better detection performance [17]. Some approaches generate synthetic foggy images for data augmentation, but these do not always capture real-world complexities [18].
@@
Moreover, domain adaptation techniques [19, 20, 21, 22, 23] have been employed to bridge the gap between foggy and clear-weather images. These methods often rely on labeled clear-weather datasets or use transfer learning, but they do not fully address the unique challenges posed by foggy conditions.
@@
Lane detection enables key functionalities such as lane departure warnings and autonomous navigation in ADAS. While existing lane detection algorithms perform well under clear weather conditions with visible lane markings, their performance drops significantly in adverse environments like fog, where reduced visibility and image degradation obscure lane clarity, thus increasing the risk of traffic accidents. Furthermore, the lack of specialized datasets tailored to foggy conditions has limited the development of effective deep learning-based solutions for these challenging scenarios.
@@
To address these limitations, we introduce the FoggyLane dataset, a new benchmark specifically designed for lane detection under foggy conditions. This dataset, comprising 1,423 annotated images across six distinct foggy road scenarios, fills a crucial gap in the current lane detection datasets. Additionally, we create foggy versions of two commonly used lane detection datasets, CULane [8] and Tusimple [24], resulting in the FoggyCULane and FoggyTusimple datasets. These datasets are essential for training and testing lane detection algorithms in foggy environments.
@@
    To enhance the utilization of structural relationships between lane instances, we introduce a Kernel Feature Fusion Module (KFFM) that learns and predicts the correlations between lane instances, improving lane detection accuracy and robustness.
@@
    To tackle the loss of edge information in foggy images, we incorporate a Low-level Edge Enhancement Module (LEEM), which improves the model’s sensitivity to fine edge details, ensuring better lane delineation even when visibility is severely impaired by fog.
@@ 2.1 Segmentation-based Lane Detection
Segmentation-based methods mainly utilize semantic segmentation or instance segmentation techniques to distinguish lane lines from other objects or backgrounds in the image, transforming lane detection into a pixel-level classification problem. LaneNet[6] employs a two-branch structure of split branches and embedded branches, the split branches output a binary lane mask to recognize lane lines, and the embedded branches assign a unique embedding to each pixel to realize the distinction of lane lines. SCNN[8] modifies the conventional layer-wise convolution approach by incorporating a slice-wise convolution technique, enabling horizontal and vertical information flow between pixels within a layer. RESA[25] introduces a spatial attention module that ensures robust lane prediction by cyclically shifting the feature map horizontally and vertically so that each pixel can acquire global information. Although these segmentation-based methods possess high accuracy and robustness, the models are bulky and slow to process due to high consumption of computational resources. HW-Transformer[26] presents an innovative lane detection network to expand the visual range around the lane and enable global information communication through intersecting features, alongside a self-attention knowledge distillation (SAKD) method to enhance performance and semantic feature learning.
@@ 2.2 Curve-based Lane Detection
Curve-based methods transform the lane detection into a regression task by predicting polynomial coefficients that fit lane lines to polynomial equations. PolyLaneNet[27] proposes an end-to-end convolutional neural network that transforms images into polynomials representing each lane’s markers, along with domain-specific lane polynomials and confidence values for every lane. Similarly, LSTR[28] utilizes a Transformer-based architecture that captures both local lane structures and broader contextual information through non-local interactions. DBNet[29] introduces NURBS curves and utilizes their geometric semantic properties to achieve local and global optimization, which improves the robustness of lane line detection and model interpretation. However, polynomial fitting is sensitive to parameter selection, and minor changes in parameters can result in significant variations in fitting results, leading to unstable lane line detection.
@@ 2.3 Anchor-based Lane Detection
Anchor-based approaches employ predetermined reference lines within the image space to generate lane predictions, subsequently calculating necessary offsets to precisely match these anchors with true lane markings. Line-CNN[30] introduces the anchor mechanism into lane detection for the first time, and innovatively proposes line anchors. LaneATT[31] effectively solves the problem of occlusion and illumination due to the lack of a line anchor, by fusing extracted anchor features with global features generated by attention module. CLRNet[16] employs a coarse-to-fine detection strategy, where lane lines are first approximated using high-level semantic cues before being refined with detailed low-level features. This hierarchical fusion of semantic information significantly improves detection precision. FLAMNet[32] employs an adaptive line anchor strategy, augmented by patch-based feature pooling and decomposed self-attention mechanisms. This design strengthens both local feature extraction and global context modeling, improving performance in handling intricate lane structures. However, due to the multiple curvature and orientation variations of lane lines in the real world, the performance of the anchor-based methods may be limited in capturing precise paths with complex shapes.
Figure 2: The route for dataset acquisition in Suzhou on the morning of October 26, 2023.[33]
Figure 3: The routes for dataset acquisition in Urumqi, Xinjiang Autonomous Region: the left illustrates the morning of December 9, 2023, while the right depicts the evening of January 2, 2024.[33]
@@ 2.4 Row-wise-based Lane Detection
Row-wise methods predict lane positions for each image row. E2E-LMD[34] introduces the row-wise-based lane detection task and develops a complete architecture that outputs lane marker positions directly. UFLD[35] utilizes global features to select the lane positions of predefined rows of an image and proposes a structural loss function, so as to ensure the continuity and smoothness of the lanes. CondLaneNet[36] utilizes conditional convolution and row-by-row prediction strategies to enhance lane recognition accuracy, and introduces the RIM module to effectively address complex lane layouts. The row-wise-based detection methods improve computational efficiency by focusing only on the row regions in the image, and possesses high detection accuracy while maintaining efficiency.
3 Data Constrution
Due to the data-driven nature of deep learning models, the quality, size, and diversity of the datasets largely determine the applicability and performance of such models. In the task of lane detection, models trained solely on clear weather data struggle to adapt to low-visibility foggy conditions. Issues such as blurred lane boundaries, faded lane colors, and reduced image quality often arise in fog, making lane detection more challenging. Therefore, training with a foggy weather lane dataset can significantly improve an autonomous driving system’s ability to handle such conditions, enhancing both detection accuracy and overall reliability. However, there is currently no publicly available foggy weather lane detection dataset. To fill this gap, we created the FoggyLane dataset for lane detection in foggy weather. In addition, we generated two supplementary datasets, FoggyCULane and FoggyTusimple, by using fog modeling techniques on existing CULane [8] and Tusimple [24] datasets to test the generalizability of our methods.
Figure 4: The different road types of FoggyLane images examples.
Dataset Frames Multi-traffic scenarios Instance-level annatation Number of foggy datasets Resolution Caltech-Lanes [37] 1.2k None 640 x 480 Tusimple [24] 6.4k None 1280 x 720 LLAMAS [38] 100k None 1276 x 717 Appolloscape [39] 110k None 3384 x 2710 BDD100k [40] 100k None 1280 x 720 CULane [8] 133k None 1640 x 590 CurveLanes [41] 150k None 1280 x 720 FoggyLane (ours) 1.423k 1.423k 1640 x 590
3.1 FoggyLane
Existing lane detection algorithms for intelligent connected vehicle ADAS systems perform well under favorable weather conditions, but their performance significantly deteriorates in adverse conditions like fog, which reduces visibility and increases the likelihood of traffic accidents. Current benchmark datasets, including CULane[8], Tusimple[24], and LLAMAS[38], lack foggy weather data. To address this, we developed FoggyLane, a dataset specifically designed for foggy conditions. Data was collected using a monocular forward-facing camera at various times and locations, including Suzhou and Urumqi, China, between October 2023 and January 2024. The dataset includes 1,086 images at various resolutions, which were resized to $1640\times 590$ for consistency.
Since foggy conditions are rare, we extended the dataset by incorporating foggy driving scenes from two YouTube videos[42], resulting in an additional 337 images. The final FoggyLane dataset contains 1,423 images, covering six road scenes: Normal, Arrow, Crowd, Curve, Night, and Crossroad. The dataset was partitioned into training, validation and testing sets, with stratification to maintain a consistent distribution of scene types across both subsets.
Comparing FoggyLane with existing datasets, such as Caltech-Lanes[37] and Tusimple[24], which focus on single traffic scenarios, and LLAMAS[38], which uses high-precision map annotations, we found that these datasets do not cover foggy weather scenarios. While ApolloScape [39] and BDD100k [40] include multiple traffic scenarios, they lack instance-level annotations. Similarly, although CULane [8] and CurveLanes [41] cover a wide range of traffic scenes, they do not include foggy conditions. Models trained on datasets limited to clear weather struggle in fog, where lane boundaries are blurred and image quality reduced. To fill this gap, FoggyLane offers a thorough solution for lane detection in foggy conditions with instance-level annotations.
Our dataset has two key advantages:
  • 1)
  • 2)
    Strict annotation standards: The dataset follows strict annotation guidelines, including marking the center of lanes, connecting lane markings at intersections, and prioritizing solid lines over grid lines, etc.[43]. These rigorous annotations can significantly improve the accuracy and consistency of lane detection in foggy weather, thus enhancing the stability and safety of autonomous driving systems and providing a solid data foundation for the reliable deployment of autonomous technology.
3.2 FoggyCULane and FoggyTusimple
Currently, numerous datasets are available for lane detection. Among them, we select two widely adopted and representative benchmarks: CULane[8] and TuSimple[24]. The CULane dataset [8] serves as an extensive and demanding benchmark, encompassing various scenarios such as highways, urban roads, and rural regions in Beijing. In contrast, the TuSimple dataset [24] primarily comprises daytime driving videos collected on American highways, featuring a variety of lighting conditions and complex traffic situations. Based on these datasets, we generate synthetic foggy versions, named FoggyCULane and FoggyTuSimple, to facilitate evaluation under adverse weather conditions.
The method for generating foggy images is inspired by the model [18], and the complete generation process is shown in Fig. 5. To be specific, the atmospheric scattering model can effectively utilize prior information to analyze the degradation mechanisms of image quality in foggy conditions. It models foggy images as a combination of scattered light and direct light based on optical principles, ultimately synthesizing a clear image into a foggy one. The mathematical modeling equation defined by Koschmieder [44] is represented as:
I(x)=J(x)t(x)+A(1t(x))I(x)=J(x)t(x)+A(1-t(x)) (1)
$t(x)=e^{-\beta d(x)}$ (2)
To calculate the corresponding $d(x)$, we use the S2R-DepthNet[45] that has good generalization performance across various applications. The S2R-DepthNet consists of three modules that decouple the structural and textural information of the input image. Then they suppress information irrelevant to depth and use structural information for depth prediction. The Structure Extraction Module (STE) breaks down the image to capture domain-independent structural details. The Depth-specific Attention (DSA) generates an attention map from the original image. Once processed through the DSA and STE Modules, as depicted in Fig. 5, the attention map and structural map are multiplied to produce a refined and all-encompassing depth-specific structure map. The Depth Prediction (DP) module is then responsible for estimating the depth.
Figure 7: The FoggyTuSimple dataset examples. (a) is the original TuSimple dataset. And (b) to (d) are the foggy images of different concentrations with $\beta$ values of 2, 4, and 8, respectively.
According to the transmittance formula Eq. (2), we use the dark channel prior[46] to calculate the average atmospheric light value $A$. In the non-sky area of an outdoor clear image, there is usually at least one very small or close to zero value in the RGB image for each pixel, which is called the dark channel of that pixel. Therefore, for any image $I$, the mathematical formula for its corresponding dark channel image $I_{\text{dark}}(x)$ can be defined as:
where $I_{\text{c}}(y)$ is the value of the c-th channel of image $I$ at pixel point $y$, and $\omega(x)$ is a small region around pixel point $x$. In the dark channel, find the top 0.1% brightest pixels and take the maximum value of their RGB channels as an approximation of the average atmospheric light value $A$. After obtaining the transmittance and the average atmospheric light value, the images of light fog, medium fog, and heavy fog can be obtained by fine-tuning the extinction coefficient $\beta$. The original images and the foggy images generated under different $\beta$ values are displayed in Figs. 6 and 7.

### v11

@@ Abstract
Lane departure accident prevention plays a critical role in enhancing road safety, and lane detection is a core technology to achieve this goal, especially under complex weather conditions. While existing lane detection algorithms perform well under favorable weather conditions, their effectiveness significantly degrades in foggy environments, which increases the risk of traffic accidents. In response to this challenge, we propose PDT-Net, a robust Prior-Guided Dynamic Feature Fusion Transformer framework designed for real-time lane detection in foggy conditions. This framework integrates three key modules: a Global Feature Fusion Module (GFFM) to capture the relationship between local and global features in foggy images, a Dynamic Feature Fusion Module (DFFM) to model the structural and positional relationships of lane instances, and a Prior-Guided Edge Enhancement Module (PEM) to recover lost edge details in foggy environments. Furthermore, we introduce the FoggyLane dataset, a real-world dataset that specifically targets lane detection in foggy conditions, along with two synthesized datasets, FoggyCULane and FoggyTusimple, to address the lack of fog-specific data for lane detection. Extensive experiments show that PDT-Net achieves state-of-the-art performance with F1-scores of 95.04% on FoggyLane, 79.85% on FoggyCULane, and 96.95% on FoggyTusimple. Moreover, with TensorRT acceleration, our method achieves a processing speed of 38.4 FPS on the NVIDIA Jetson AGX Orin, confirming its real-time capability and robustness in challenging foggy environments. By improving the precision of lane detection, our framework can contribute to active safety warning systems, helping to prevent accidents in foggy conditions.
@@ 1 Introduction
Road traffic crashes continue to be a major global safety concern, with lane departure being a significant contributor. According to the National Highway Traffic Safety Administration (NHTSA), over 40% of fatal accidents in 2023 were attributed to lane departure(NHTSA2025), primarily caused by adverse weather conditions, degraded road markings, and driver errors. These factors expose critical gaps in vehicle perception, highlighting the need for more robust lane detection systems to address this issue.
Computer vision techniques, particularly those focused on real-time risk assessment and accident prediction, play a pivotal role in enhancing active safety warning systems(AAP4). Lane detection, as a core perceptual capability, enables Advanced Driver Assistance Systems (ADAS) to trigger active safety warnings such as lane departure warning, which are essential for accident prevention.
@@
Under good weather and clear lane markings, current lane detection algorithms perform well. However, in complex road environments and poor weather, such as those with poor weather conditions like fog, lane markings can become blurred, image quality may decrease, and global environmental information is often obscured, which makes lane detection more challenging. In these conditions, existing algorithms often struggle and may fail, leading to an increased risk of traffic accidents with more severe consequences. Ensuring safe driving in these conditions calls for the development of high-precision, highly robust lane detection methods optimized for foggy environments. As shown in Fig. 1, The in-vehicle camera captures real-time road images, which are then preprocessed by edge devices and sent to the lane detection module. The detected lane information is used for active safety alerts. In the event of lane departure, the system proactively issues warnings(AAP7) through steering wheel vibration, voice notifications, and seat vibration, helping to prevent accidents(AAP8).
@@
Earlier lane detection algorithms(son2015real; hou2016efficient) primarily relied on traditional computer vision methods, such as feature-based approaches that detect vanishing points, use color features, and apply edge detection or line detection techniques like Hough transforms. These methods are efficient for real-time applications but require manual parameter tuning and often struggle in challenging scenarios like worn or occluded lane markings, which may lead to serious collision accidents.
Figure 1: Lane detection in foggy scenarios for ADAS.(Pexels) Inaccurate lane detection may lead to lane departure, increasing the risk of collision. Accurate lane detection provides essential technical support for the active safety functions of vehicles. Once the vehicle deviates from its intended lane, the ADAS can promptly issue an active safety warning to alert the driver, enabling timely corrective actions and ultimately preventing the occurrence of traffic accidents.
@@
With the rise of deep learning, many modern lane detection methods (neven2018towards; yu2018deep; pan2018spatial; abualsaud2021laneaf) leverage CNNs, which automatically extract features from raw data and generalize better in complex road environments. While deep learning provides higher robustness, many models still rely on pixel-by-pixel classification, which is computationally expensive and unnecessary for lane detection. Moreover, lane detection doesn’t always require dense pixel segmentation, which can be inefficient.
@@
Additionally, some approaches have incorporated anchor-based methods from object detection to improve efficiency and accuracy, particularly for complex or occluded lanes. These methods preset anchors at different starting points and slopes to optimize lane shape and position prediction (AAP1; redmon2016you; farhadi2018yolov3; chen2019pointlanenet; zheng2022clrnet; AAP5). However, these techniques still face challenges in handling foggy conditions, where image degradation severely impacts lane visibility. Traditional image enhancement methods, like dehazing, are commonly used to improve clarity but do not always result in better detection performance (huang2020dsnet). Some approaches generate synthetic foggy images for data augmentation, but these do not always capture real-world complexities (nie2022sensor).
@@
Moreover, domain adaptation techniques (chen2018domain; zhang2021domain; hnewa2021multiscale; li2023domain) have been employed to bridge the gap between foggy and clear-weather images. These methods often rely on labeled clear-weather datasets or use transfer learning, but they do not fully address the unique challenges posed by foggy conditions.
@@
Lane detection enables key functionalities such as lane departure warnings and autonomous navigation in ADAS, demonstrating strong risk perception capabilities(AAP3; AAP6). While existing lane detection algorithms perform well under clear weather conditions with visible lane markings, their performance drops significantly in adverse environments like fog, where reduced visibility and image degradation obscure lane clarity, thus increasing the risk of traffic accidents. Furthermore, the lack of specialized datasets tailored to foggy conditions has limited the development of effective deep learning-based solutions for these challenging scenarios.
@@
To address these limitations, we introduce the FoggyLane dataset, a new benchmark specifically designed for lane detection under foggy conditions. This dataset, comprising 1,423 annotated images across six distinct foggy road scenarios, fills a crucial gap in the current lane detection datasets. Additionally, we create foggy versions of two commonly used lane detection datasets, CULane (pan2018spatial) and Tusimple (Tusimple), resulting in the FoggyCULane and FoggyTusimple datasets. These datasets are essential for training and testing lane detection algorithms in foggy environments.
@@
    To enhance the utilization of structural relationships between lane instances, we introduce a Dynamic Feature Fusion Module (DFFM) that learns and predicts the correlations between lane instances, improving lane detection accuracy and robustness.
@@
    To tackle the loss of edge information in foggy images, we incorporate a Prior-Guided Edge Enhancement Module (PEM), which enhances the model’s sensitivity to fine edge details. This ensures better lane delineation, even when visibility is severely impaired by fog, and significantly improves the reliability of lane departure warnings.
@@ 2.1 Segmentation-based Lane Detection
Segmentation-based methods mainly utilize semantic segmentation or instance segmentation techniques to distinguish lane lines from other objects or backgrounds in the image, transforming lane detection into a pixel-level classification problem. LaneNet(neven2018towards) employs a two-branch structure of split branches and embedded branches, the split branches output a binary lane mask to recognize lane lines, and the embedded branches assign a unique embedding to each pixel to realize the distinction of lane lines. SCNN(pan2018spatial) modifies the conventional layer-wise convolution approach by incorporating a slice-wise convolution technique, enabling horizontal and vertical information flow between pixels within a layer. RESA(zheng2021resa) introduces a spatial attention module that ensures robust lane prediction by cyclically shifting the feature map horizontally and vertically so that each pixel can acquire global information. Although these segmentation-based methods possess high accuracy and robustness, the models are bulky and slow to process due to high consumption of computational resources. HW-Transformer(tits3) presents an innovative lane detection network to expand the visual range around the lane and enable global information communication through intersecting features, alongside a self-attention knowledge distillation (SAKD) method to enhance performance and semantic feature learning.
@@ 2.2 Curve-based Lane Detection
Curve-based methods transform the lane detection into a regression task by predicting polynomial coefficients that fit lane lines to polynomial equations. PolyLaneNet(tabelini2021polylanenet) proposes an end-to-end convolutional neural network that transforms images into polynomials representing each lane’s markers, along with domain-specific lane polynomials and confidence values for every lane. Similarly, LSTR(liu2021end) utilizes a Transformer-based architecture that captures both local lane structures and broader contextual information through non-local interactions. DBNet(dai2024dbnet) introduces NURBS curves and utilizes their geometric semantic properties to achieve local and global optimization, which improves the robustness of lane line detection and model interpretation. However, polynomial fitting is sensitive to parameter selection, and minor changes in parameters can result in significant variations in fitting results, leading to unstable lane line detection.
@@ 2.3 Anchor-based Lane Detection
Anchor-based approaches employ predetermined reference lines within the image space to generate lane predictions, subsequently calculating necessary offsets to precisely match these anchors with true lane markings. Line-CNN(li2019line) introduces the anchor mechanism into lane detection for the first time, and innovatively proposes line anchors. LaneATT(tabelini2021keep) effectively solves the problem of occlusion and illumination due to the lack of a line anchor, by fusing extracted anchor features with global features generated by attention module. CLRNet(zheng2022clrnet) employs a coarse-to-fine detection strategy, where lane lines are first approximated using high-level semantic cues before being refined with detailed low-level features. This hierarchical fusion of semantic information significantly improves detection precision. FLAMNet(ran2023flamnet) employs an adaptive line anchor strategy, augmented by patch-based feature pooling and decomposed self-attention mechanisms. This design strengthens both local feature extraction and global context modeling, improving performance in handling intricate lane structures. However, due to the multiple curvature and orientation variations of lane lines in the real world, the performance of the anchor-based methods may be limited in capturing precise paths with complex shapes.
@@ 2.4 Row-wise-based Lane Detection
Row-wise methods predict lane positions for each image row. E2E-LMD(yoo2020end) introduces the row-wise-based lane detection task and develops a complete architecture that outputs lane marker positions directly. UFLD(qin2020ultra) utilizes global features to select the lane positions of predefined rows of an image and proposes a structural loss function, so as to ensure the continuity and smoothness of the lanes. CondLaneNet(liu2021condlanenet) utilizes conditional convolution and row-by-row prediction strategies to enhance lane recognition accuracy, and introduces the RIM module to effectively address complex lane layouts. The row-wise-based detection methods improve computational efficiency by focusing only on the row regions in the image, and possesses high detection accuracy while maintaining efficiency.
Figure 2: Overview of the proposed network.
Figure 3: Backbone Network with Global Feature Fusion Module(GFFM).